NASA ISS On-Orbit Status 25 May 2015

©NASA

NASA ISS On-Orbit Status 25 May 2015.

Today: Fluid Shifts Before, During and After Prolonged Space Flight and Their Association with Intracranial Pressure and Visual Impairment (Fluid Shifts): Kelly and Padalka performed a system test of the Ultrasound in the Service Module.

The crew configured the ultrasound, laptop, and video, and with assistance from the ground remote guider, then performed an ultrasound scan. This test of capability is to verify capability in advance of the Fluid Shift ultrasound operations when the crew uses the Lower Body Negative Pressure (Chibis) suit next week. This week, the crew will be performing their baseline data collections in the United States On-orbit Segment (USOS). Fluid Shifts is a joint NASA-Russian experiment that investigates the causes for severe and lasting physical changes to astronaut's eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is investigated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and eye damage.

Marangoni Deformation 30 (Dynamic Surf-3): Cristoforetti set up the Fluid Physics Experiment Facility (FPEF) Dynamic Surf-3 core and body. She then configured the Dynamic Surf-3 cassette, installed it into the core and body. The assembly was then installed into the FPEF. Cristoforetti completed the Dynamic Surf-3 setup activities on Wednesday, with the ground checkout activities later that night. The Dynamic Surf investigation is part of a series of JAXA experiments that Marangoni convection driven by the presence of surface tension gradient as produced by a temperature difference at a liquid/gas interface. Fluid convection observations of a silicone oil liquid bridge that is generated by heating the one disc higher than the other within the Fluid Physics Experiment Facility (FPEF). By observing and understanding how such fluids move researchers can learn about how heat is transferred in microgravity, and ultimately drive the design and development of more efficient fluid flow based systems and devices. The objective of scientific research on Marangoni convection utilizing microgravity is to make clear the flow transition phenomena from steady to oscillatory, chaotic, and finally turbulent flows. Therefore, it is important to understand an underlying principal of Marangoni convection. The findings and knowledge obtained through the space experiment is applied to industrial processes, as well as fluid physics.

Handhold Experiment Platform (HXP) Installation: This week the crew and ground controllers will be installing the HXP on the JEM Exposed Facility (JEF). Today, ground controllers used the JEM Remote Manipulator System (JEMRMS) to retrieve the Small Fine Arm (SFA). On Tuesday, the JEM Airlock (JEMAL) outer hatch will be opened and slide table extended for the SFA to grapple the HXP and install on one of the handrails of the JEF. The HXP is composed of several exposure experiments which will remain attached to the JEF for approximately one year, collecting data on material interactions to long duration space exposure.

Extravehicular Mobility Unit (EMU) Loop Scrub: Virts configured EMU suits 3003 and 3010 for today's activities. He then successfully initiated ionic and particulate filtration (scrubbing) on both suits. Later during biocidal maintenance (iodination), EMU 3010 was unable to establish sufficient cooling water flow to complete the process. Iodination maintenance was successfully performed on EMU 3003. A water sample was obtained to determine the effectiveness of the filtering. A portion of this water sample will used for a conductivity test and the remaining water will be sent to the ground for chemical analysis. Ground Teams are currently planning to perform a Fan Pump Separator change out on EMU 3010 during the first week of June.

Permanent Multipurpose Module (PMM) Preparation: In preparation for the upcoming PMM relocation, Kelly has secured PMM Rack hardware and stowage. In addition, he gathered vestibule outfitting hardware for Node 1 Nadir and Node 3 Forward ports later in the day. PMM relocation from Node 1 Nadir to Node 3 Forward is planned for Wednesday, May 27th starting at 4:30am CDT.

Today's Planned Activities
All activities were completed unless otherwise noted.

MORZE. Bioimpedance measurements with use of SPRUT-2 device. Tag-up with the specialists
CORRECTSIA. Food and liquids (medical preparations) ingestion to be registered in the log.
[ ВКС] laptops antivirus base update
ARED - wire ropes lubrication
JEMRMS bus monitor cable installation
EMU suits backpack change out
WRS - processing tank fill out from EDV container
PUMA device routine test check out
JEMRMS manipulator bus monitor activation and start up [Aborted]
EMU 3003 suit and EMU 3010 cooling loop cleaning initiation
MORZE. Psychophysiological Evaluation обследование :Tsentrovka f tests SENSOR
WRS - Water Samples Analysis
Final operations prior to PMM hatches closure
KARDIOVEKTOR. Experiment run. Tag-up with the specialists if required
FS- Set up of hardware in SM
FLUID SHIFTS. Set up and connection of hardware for TV Footage (Ku+S-band)
JEM - Start of gas mixture delivery.
FLUID SHIFTS. Communication assets configuration for experiment run
JEMRMS Arm monitoring - termination [Aborted]
Life on the Station. Photo and video coverage
EMU-H2O - conductivity test
FLUID SHIFTS.USDN Connection in SM
FS-installation of hardware in SM
Node3 forward hatch area clearing
Water sampling after EMU suit cooling loop scrubbing
MD30 - Task Familiarization
LBNP Exercise (PRELIMINARY) - Assistance
FS- Ultrasound survey in SM
LBNP Exercise (PRELIMINARY) Tag-up with the specialists
EMU-H2O - conductivity test
XF305- Installation of video camera
MD30- Preparation for the experiment Run
FS- Hardware Disassemble in SM with photo footage
Harmful contaminants measurement with the use of CMS в СМ
Installation of shunting plugs on [СтМ ТКГ №425(АО)]. Re-stowage of the bags with the plugs
EMU Suit Cooling Loop scrub - Iodination
FLUID SHIFTS. Disconnecting Ultrasound P\L
FLUID SHIFTS. КСПЭ hardware Deactivation and Closing Application on SSC ЦП
FS- Transfer of hardware for stowage from SM
TOCA - Data Recording
TOCA - WWB replacement
FLUID SHIFTS. Return to nominal Comm Configuration
CORRECTSIA. Food and liquids (medical preparations) ingestion to be registered in the log
Installation of inserts for easy opening of SM interior panel locks (panel 437) (Kit with structural elements (004890R, ФГБ1ПГО_4_427_1, CTB No.1112 (004106J)
PROBOY. RSE1 laptop Ops. Tag-up with the specialists
PROBOY. Penetration Simulator Ops. Tag-up with the specialists
MD30- Installation of the cassette into Equipment
Hardware Gathering for PMM Vestibule Outfitting
EMU Suits Configuration for cooling loop cleaning
Filling (Separation) of EDV (KOV) (Separate No.1162 (00063575R, ФГБ1ПГО_1_110) for Elektron For separation use additional EDV (KOV) No.1104 (00063568R, ФГБ1ПГО_4_404_1)
EMU - Drying
PROBOY. Copying and downlink of the data to the Ground
СОЖ Maintenance
Node2 hardware re-stowage
MORZE. Psychophysiological Evaluation: SUPOS test
EMU Suit Configuration tear down
MD30- Installation of Equipment
MATRESHKA-R. Collection and measurements of Bubble-dosimeter sensors. Tag-up with the specialists if required
Combustion products analyzer routine maintenance.
ВШТВ optical device functionality check
IMS data edit
Physical Exercise data downlink
Preparation for anti-virus scanning on [ВКС] laptops
CORRECTSIA. Food and liquids (medical preparations) ingestion to be registered in the log
Fine Motor Skills First - Reading the Reminder
Preparation of reports for Roscosmos site
ECON-M. Observations and Photography
Completed Task List Items

None
Ground Activities
All activities were completed unless otherwise noted.

EMU Loop Scrub Support
Maneuver JEMRMS SFA to checkout position
Three-Day Look Ahead:

Tuesday, 05/26: Node 3 CBCS Install, N1 vestibule config + CPA Install, N1 Vestibule depress
Wednesday, 05/27: PMM Relocate and Install, Fluid Shifts
Thursday, 05/28: PMM Vestibule outfitting, Node 3 Fwd Starboard IMV Valve Install, Fluid Shifts
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - On
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Shutdown
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Standby
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.